y=-x+5
y=2x-4
This is the problem I have to do. Can you use any method for solving systems of equation? Can you also explain the problem?

Answer by jim_thompson5910(35256)

Since y is both -x+5 and 2x-4, we can say

y = -x+5

2x - 4 = -x + 5

Solve that last equation for x. Then use the solution you found for x to find the solution for y.